OpenStudy (anonymous):

if f(x)=2x+1 and g(x)=x^2-7 find (fog)(2)

OpenStudy (anonymous):

@iGreen please can you help

OpenStudy (igreen):

I think you multiply \((2x + 1)(x^2-7)\)., then multiply it by 2..

OpenStudy (anonymous):

thats what i was thinking but im not 100% on it

OpenStudy (igreen):

Can you multiply \((2x + 1)(x^2 - 7)\)?

OpenStudy (anonymous):

yep give me a second

OpenStudy (anonymous):

I;ve gotten -5

OpenStudy (anonymous):

is that what you got too?

OpenStudy (anonymous):

2(x^2-7)+1 2(2^2-7)+1 2(4-7)+1 2(-3)+1 -6+1=-5 answer: -5

OpenStudy (igreen):

....... How do you get -5?

OpenStudy (igreen):

Um..no I think that's wrong.

OpenStudy (igreen):

\((2x + 1)(x^2 - 7) = (2x^3 + x^2 - 14x - 7)(2) = 4x^3 + 2x^2 - 28x - 14\)

OpenStudy (igreen):

Is that one of your answer choices?

OpenStudy (anonymous):

it was an essay question unfortunately

OpenStudy (anonymous):

i'll put your one down too just in case :P thanks for your help @iGreen

OpenStudy (igreen):

No problem.